Three people are selected at random from four females and eight
males. Find the probability of the following. (Round your answers
to three decimal places.)
(a) At least one is a male.
(b) At most two are male.

Answer by Edwin McCravy(20060)

Three people are selected at random from four females and eight
males. Find the probability of the following. (Round your answers
to three decimal places.)
(a) At least one is a male.
The complement event is the event that 3 females are selected.

P(complement event) = 

(4 females CHOOSE 3)/(12 people CHOOSE 3) = C(4,3)/C(12,3) = ____?

1 - P(complement event) = 1 - ____ = ____?

-------------------------------------------------------------
(b) At most two are male.
The complement event is the event that 3 males are selected.

P(complement event) = 

(8 males CHOOSE 3)/(12 people CHOOSE 3) = C(8,3)/C(12,3) = ____?

1 - P(complement event) = 1 - ____ = ____?

Edwin